### v1

@@ Abstract
Traditional wave forecasting models, although based on energy conservation equations, are computationally expensive. On the other hand, existing deep learning geophysical fluid models, while computationally efficient, often suffer from issues such as energy dissipation in long-term forecasts. This paper proposes a novel energy-balanced deep learning wave forecasting model called WaveCastNet (WCN). By incorporating wind fields at the current, previous, and future time steps, as well as wave fields at the current and previous time steps as input variables, WCN maintains energy balance within the model. Furthermore, the model employs adaptive Fourier operators as its core components and designs a masked loss function to better handle the impact of land-sea boundaries. A series of experiments on the ERA5 dataset demonstrate that WCN can achieve short-term forecast accuracy comparable to traditional models while exhibiting an understanding of the wave generation process. In comparative experiments under both normal and extreme conditions, WCN consistently outperforms the widely used WaveWatch III model in the industry. Even after long-term forecasting, WCN maintains a stable and energy-rich state. By further constructing a simple meteorological model, WCN-wind, which considers energy balance, this paper confirms the importance of energy constraints for improving the long-term forecast performance of deep learning meteorological models. This finding provides new ideas for future research on deep learning geophysical fluid models.
@@ 2 Data and Methods
where represents the significant wave height at time , and are the meridional and zonal wind speeds at 10 m above the sea surface, respectively, and represents the model’s mapping function. We name the model WaveCastNet (WCN).
@@ 3.2 Performance Comparison against Traditional Forecasting Models
In wave forecasting, WW3 model is one of the most powerful traditional forecasting models. As mentioned earlier, WW3 is a third-generation wave model based on the energy balance equation. For the comparison data, we used WW3 Global Wave Model data from the University of Hawaii, which is driven by NOAA/NCEP’s Global Forecast System (GFS) winds. This dataset provides hourly wave forecasting data at 0.5-degree spatial resolution. When comparing with the WCN model, our model’s initial fields were derived from ERA5 data, and the wind fields used for forcing were the same as those used in the WW3 model. As model errors accumulate over time, shorter forecast lead times generally result in higher model accuracy. The University of Hawaii’s WW3 model updates its forecasts once a day and only retains the forecast data for the next 24 hours, as this period has the smallest forecast error and the highest reliability. Therefore, when comparing with the WCN model, we selected this ’24-hour short-term forecast’ result from the WW3 model. Since the University of Hawaii’s model started forecasting in 2023, we selected the main impact periods of the strongest typhoon (Mawar) and the strongest hurricane (Lee) in 2023 to validate the model’s forecasting performance. Additionally, we randomly selected one day from each month, excluding these two periods, and compared the model’s performance based on the average results. As the ERA5 reanalysis data for December 2023 was not fully updated at the time of writing this article, we only sampled from the first 11 months.
@@
During the first 7 time steps (1.75 days) of the forecasting process, the MAE of the WCN model was 0.14, which was lower than the MAE of the WW3 model (0.30). In terms of ACC, the minimum value for the WCN model was 0.99, higher than the 0.98 of the WW3 model. Regarding RMSE, the maximum value for the WCN model was 0.02, lower than the 0.03 of the WW3 model. Considering the evaluation curves, we can conclude that the short-term forecasting performance of the WCN model under normal conditions is better than that of the WW3 model.
@@
Typhoon Mawar (2023) was the strongest typhoon in the Pacific region in 2023, forming on May 19 and dissipating on June 11. It reached its peak intensity of 68 m/s on May 26. Therefore, we selected May 25 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on May 24 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the WCN model and the WW3 model. Some of the results are shown in Figure 5.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 120°E to 162°E, latitude 0°N to 30°N), the MAE of the WCN model (0.16 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the WCN model (0.98) was higher than that of the WW3 model (0.97). Regarding RMSE, the maximum value for the WCN model (0.07 m) was lower than that of the WW3 model (0.09 m). Considering the evaluation curves, we can conclude that the WCN model performed better than the WW3 model in forecasting the Mawar region.
Hurricane Lee (2023) was the strongest hurricane in the Atlantic region in 2023, forming on September 5 and dissipating on September 18. It rapidly intensified into a Category 5 hurricane on September 7, with its wind speed increasing by 85 mph (140 km/h) within 24 hours, becoming the third-fastest rapid intensification on record in the Atlantic. Rapid intensification refers to a significant increase in a tropical cyclone’s maximum sustained winds within a short period, typically 24 hours or less. We selected September 6 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on September 5 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the WCN model and the WW3 model. Some of the results are shown in Figure 6.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 72°W to 30°W, latitude 0°N to 30°N), the MAE of the WCN model (0.17 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the WCN model (0.79) was slightly lower than that of the WW3 model (0.82). Regarding RMSE, the maximum value for the WCN model (0.03 m) was similar to that of the WW3 model (0.03 m). However, if we focus on the first 4 time steps (before the WW3 model updates its forecasting results), the minimum ACC value for the WCN model (0.98) was higher than that of the WW3 model (0.96), and the maximum RMSE value for the WCN model (0.02 m) was lower than that of the WW3 model (0.03 m). Based on these results, we can conclude that the short-term forecasting performance of the WCN model is comparable to that of traditional forecasting models.
@@ 4 Discussion
The reason for this difference is that when constructing the wave model, we considered wind energy and previous wave energy as inputs, thereby maintaining energy balance within the model. In contrast, current deep learning meteorological models do not explicitly consider the issue of energy conservation. To further demonstrate the importance of energy balance for improving model performance, we attempted to construct a simple meteorological model, WaveCastNet-wind (WCN-wind). This model takes the 10-m wind field at the sea surface as the forecasting object and selects factors that have a significant impact on wind field energy changes as input variables, including the wind fields at the current and previous time steps, as well as solar radiation intensity at the current, previous, and future time steps. Specifically, the input-output relationship of the model can be represented by the following function:
@@
We used the same training data and methods as WCN to train WCN-wind and compared its long-term forecasting results with FourCastNet, which is also based on adaptive Fourier operators. As shown in Figure 11, at the 75th forecasting time step, FourCastNet’s forecasting results have already exhibited significant energy dissipation phenomena, with some areas even experiencing numerical instability issues. In contrast, the forecasting results of WCN-wind always maintain a reasonable energy level without any signs of collapse. This result proves that by introducing energy balance constraints into the model, the long-term forecasting performance of deep learning meteorological models can be effectively improved.
@@ 5 Conclusion
In this study, we developed a deep learning-based wave forecasting model WaveCastNet (WCN) and compared its performance with a traditional physical model, WaveWatch III (WW3). Our results demonstrate that the WCN model can effectively learn the intrinsic relationships between wind and waves, capturing the complex physical processes involved in wave generation, propagation, and dissipation.
Through a series of experiments, we found that the WCN model exhibits superior performance compared to traditional models in both normal and extreme conditions. In the comparison with the WW3 model, WCN showed lower mean absolute errors, higher anomaly correlation coefficients, and lower root mean square errors, indicating its ability to provide more accurate wave forecasts. Furthermore, the WCN model demonstrated a convergence process similar to that of physical models, even under the influence of extreme wind fields, such as those associated with typhoons and hurricanes.
Our analysis of the energy density in the model results revealed that the deep learning process not only learns the statistical relationships between variables but also captures the underlying physical mechanisms. By comparing the wind energy density and wave energy density curves of different models, we found that the WCN model maintains a better balance between energy input and dissipation, closely resembling the actual observation results. In contrast, some autoregressive models, such as FUXI and FourCastNet, exhibited significant energy dissipation over time, highlighting the importance of incorporating energy input in long-term forecasting.
Moreover, our study sheds light on the potential limitations of current autoregressive models in meteorological and oceanographic applications. These models often rely solely on the previous time step’s forecast results, neglecting other crucial inputs, such as solar radiation, which drives atmospheric and oceanic circulations. By incorporating energy input terms, as demonstrated in the WCN-wind model, we showed that deep learning models can maintain more realistic wind field characteristics over extended simulation periods.
In conclusion, our findings suggest that deep learning models, such as WCN, have the potential to revolutionize wave forecasting and other geophysical applications. By learning the intrinsic physical processes and incorporating essential energy input terms, these models can provide accurate and reliable long-term predictions, outperforming traditional physical models and existing deep learning approaches. Future research should focus on further improving the representation of physical processes in deep learning models and exploring their applicability to a wider range of geophysical phenomena.
@@ Acknowledgments
We extend our special thanks to the authors of FourCastNet for their inspiring work. The architecture of our WaveCastNet model is built upon the framework of FourCastNet, and we are grateful for their innovative contributions to the field of deep learning-based weather forecasting.

### v2

@@ Abstract
Traditional wave forecasting models, although based on energy conservation equations, are computationally expensive. On the other hand, existing deep learning geophysical fluid models, while computationally efficient, often suffer from issues such as energy dissipation in long-term forecasts. This paper proposes a novel energy-balanced deep learning wave forecasting model called OceanCastNet (OCN). By incorporating wind fields at the current, previous, and future time steps, as well as wave fields at the current and previous time steps as input variables, OCN maintains energy balance within the model. Furthermore, the model employs adaptive Fourier operators as its core components and designs a masked loss function to better handle the impact of land-sea boundaries. A series of experiments on the ERA5 dataset demonstrate that OCN can achieve short-term forecast accuracy comparable to traditional models while exhibiting an understanding of the wave generation process. In comparative experiments under both normal and extreme conditions, OCN consistently outperforms the widely used WaveWatch III model in the industry. Even after long-term forecasting, OCN maintains a stable and energy-rich state. By further constructing a simple meteorological model, OCN-wind, which considers energy balance, this paper confirms the importance of energy constraints for improving the long-term forecast performance of deep learning meteorological models. This finding provides new ideas for future research on deep learning geophysical fluid models.
@@ 2 Data and Methods
where represents the significant wave height at time , and are the meridional and zonal wind speeds at 10 m above the sea surface, respectively, and represents the model’s mapping function. We name the model OceanCastNet (OCN).
@@ 3.2 Performance Comparison against Traditional Forecasting Models
In wave forecasting, WW3 model is one of the most powerful traditional forecasting models. As mentioned earlier, WW3 is a third-generation wave model based on the energy balance equation. For the comparison data, we used WW3 Global Wave Model data from the University of Hawaii, which is driven by NOAA/NCEP’s Global Forecast System (GFS) winds. This dataset provides hourly wave forecasting data at 0.5-degree spatial resolution. When comparing with the OCN model, our model’s initial fields were derived from ERA5 data, and the wind fields used for forcing were the same as those used in the WW3 model. As model errors accumulate over time, shorter forecast lead times generally result in higher model accuracy. The University of Hawaii’s WW3 model updates its forecasts once a day and only retains the forecast data for the next 24 hours, as this period has the smallest forecast error and the highest reliability. Therefore, when comparing with the OCN model, we selected this ’24-hour short-term forecast’ result from the WW3 model. Since the University of Hawaii’s model started forecasting in 2023, we selected the main impact periods of the strongest typhoon (Mawar) and the strongest hurricane (Lee) in 2023 to validate the model’s forecasting performance. Additionally, we randomly selected one day from each month, excluding these two periods, and compared the model’s performance based on the average results. As the ERA5 reanalysis data for December 2023 was not fully updated at the time of writing this article, we only sampled from the first 11 months.
@@
During the first 7 time steps (1.75 days) of the forecasting process, the MAE of the OCN model was 0.14, which was lower than the MAE of the WW3 model (0.30). In terms of ACC, the minimum value for the OCN model was 0.99, higher than the 0.98 of the WW3 model. Regarding RMSE, the maximum value for the OCN model was 0.02, lower than the 0.03 of the WW3 model. Considering the evaluation curves, we can conclude that the short-term forecasting performance of the OCN model under normal conditions is better than that of the WW3 model.
@@
Typhoon Mawar (2023) was the strongest typhoon in the Pacific region in 2023, forming on May 19 and dissipating on June 11. It reached its peak intensity of 68 m/s on May 26. Therefore, we selected May 25 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on May 24 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the OCN model and the WW3 model. Some of the results are shown in Figure 5.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 120°E to 162°E, latitude 0°N to 30°N), the MAE of the OCN model (0.16 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.98) was higher than that of the WW3 model (0.97). Regarding RMSE, the maximum value for the OCN model (0.07 m) was lower than that of the WW3 model (0.09 m). Considering the evaluation curves, we can conclude that the OCN model performed better than the WW3 model in forecasting the Mawar region.
Hurricane Lee (2023) was the strongest hurricane in the Atlantic region in 2023, forming on September 5 and dissipating on September 18. It rapidly intensified into a Category 5 hurricane on September 7, with its wind speed increasing by 85 mph (140 km/h) within 24 hours, becoming the third-fastest rapid intensification on record in the Atlantic. Rapid intensification refers to a significant increase in a tropical cyclone’s maximum sustained winds within a short period, typically 24 hours or less. We selected September 6 as the starting point for forecasting, using the ERA5 wave and wind fields at 12:00 and 18:00 on September 5 as the initial conditions, and the GFS data as the wind forcing. We conducted a 3-day forecast and compared the differences between the OCN model and the WW3 model. Some of the results are shown in Figure 6.
During the 11-time-step (2.75 days) forecasting period, within the forecasting area (longitude 72°W to 30°W, latitude 0°N to 30°N), the MAE of the OCN model (0.17 m) was lower than that of the WW3 model (0.26 m). In terms of ACC, the minimum value for the OCN model (0.79) was slightly lower than that of the WW3 model (0.82). Regarding RMSE, the maximum value for the OCN model (0.03 m) was similar to that of the WW3 model (0.03 m). However, if we focus on the first 4 time steps (before the WW3 model updates its forecasting results), the minimum ACC value for the OCN model (0.98) was higher than that of the WW3 model (0.96), and the maximum RMSE value for the OCN model (0.02 m) was lower than that of the WW3 model (0.03 m). Based on these results, we can conclude that the short-term forecasting performance of the OCN model is comparable to that of traditional forecasting models.
@@ 4 Discussion
The reason for this difference is that when constructing the wave model, we considered wind energy and previous wave energy as inputs, thereby maintaining energy balance within the model. In contrast, current deep learning meteorological models do not explicitly consider the issue of energy conservation. To further demonstrate the importance of energy balance for improving model performance, we attempted to construct a simple meteorological model, OceanCastNet-wind (OCN-wind). This model takes the 10-m wind field at the sea surface as the forecasting object and selects factors that have a significant impact on wind field energy changes as input variables, including the wind fields at the current and previous time steps, as well as solar radiation intensity at the current, previous, and future time steps. Specifically, the input-output relationship of the model can be represented by the following function:
@@
We used the same training data and methods as OCN to train OCN-wind and compared its long-term forecasting results with FourCastNet, which is also based on adaptive Fourier operators. As shown in Figure 11, at the 75th forecasting time step, FourCastNet’s forecasting results have already exhibited significant energy dissipation phenomena, with some areas even experiencing numerical instability issues. In contrast, the forecasting results of OCN-wind always maintain a reasonable energy level without any signs of collapse. This result proves that by introducing energy balance constraints into the model, the long-term forecasting performance of deep learning meteorological models can be effectively improved.
@@ 5 Conclusion
In this study, we developed a deep learning-based wave forecasting model OceanCastNet (OCN) and compared its performance with a traditional physical model, WaveWatch III (WW3). Our results demonstrate that the OCN model can effectively learn the intrinsic relationships between wind and waves, capturing the complex physical processes involved in wave generation, propagation, and dissipation.
Through a series of experiments, we found that the OCN model exhibits superior performance compared to traditional models in both normal and extreme conditions. In the comparison with the WW3 model, OCN showed lower mean absolute errors, higher anomaly correlation coefficients, and lower root mean square errors, indicating its ability to provide more accurate wave forecasts. Furthermore, the OCN model demonstrated a convergence process similar to that of physical models, even under the influence of extreme wind fields, such as those associated with typhoons and hurricanes.
Our analysis of the energy density in the model results revealed that the deep learning process not only learns the statistical relationships between variables but also captures the underlying physical mechanisms. By comparing the wind energy density and wave energy density curves of different models, we found that the OCN model maintains a better balance between energy input and dissipation, closely resembling the actual observation results. In contrast, some autoregressive models, such as FUXI and FourCastNet, exhibited significant energy dissipation over time, highlighting the importance of incorporating energy input in long-term forecasting.
Moreover, our study sheds light on the potential limitations of current autoregressive models in meteorological and oceanographic applications. These models often rely solely on the previous time step’s forecast results, neglecting other crucial inputs, such as solar radiation, which drives atmospheric and oceanic circulations. By incorporating energy input terms, as demonstrated in the OCN-wind model, we showed that deep learning models can maintain more realistic wind field characteristics over extended simulation periods.
In conclusion, our findings suggest that deep learning models, such as OCN, have the potential to revolutionize wave forecasting and other geophysical applications. By learning the intrinsic physical processes and incorporating essential energy input terms, these models can provide accurate and reliable long-term predictions, outperforming traditional physical models and existing deep learning approaches. Future research should focus on further improving the representation of physical processes in deep learning models and exploring their applicability to a wider range of geophysical phenomena.
@@ Acknowledgments
We extend our special thanks to the authors of FourCastNet for their inspiring work. The architecture of our OceanCastNet model is built upon the framework of FourCastNet, and we are grateful for their innovative contributions to the field of deep learning-based weather forecasting.